### v1

# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

Abstract
Training more counselors, from clinical students to peer supporters, can help meet the demand for accessible mental health support; however, current training approaches remain resource-intensive and difficult to scale effectively. Large Language Models (LLMs) offer promising solutions for growing counseling skills training through simulated practice and automated feedback. Despite successes in aligning LLMs with expert-counselor annotations, we do not know whether LLM-based counseling training tools—such as AI patients that simulate real-world challenges and generative AI feedback with suggested alternatives and rationales—actually lead to improvements in novice counselor skill development. We develop CARE, an LLM-simulated practice and feedback system, and randomize 94 novice counselors to practice using an AI patient, either alone or with AI feedback, measuring changes in their behavioral performance, self-assessments, and qualitative learning takeaways. Our results show the practice-and-feedback group improved in their use of reflections and questions (d=0.32-0.39, p$<$0.05). In contrast, the group that practiced with an AI patient alone did not show improvements, and in the case of empathy, actually had worse uses across time (d=--0.52, p=0.001) and when compared against the practice-and-feedback group (d=0.72, p=0.001). Participants’ qualitative self-reflections revealed key differences: the practice-and-feedback group adopted a client-centered approach involving listening to and validating feelings, while the practice-alone group remained solution-oriented but delayed offering suggestions until gathering more information. Overall, these results suggest that LLM-based training systems can promote effective skill development, but that combining both simulated practice and structured feedback is critical.
In 2023, 22.8% of U.S. adults (approximately 58.7 million people) experienced a mental illness [1]. Unfortunately, the ability to provide people with effective mental health care is severely limited by shortages of high-quality mental health providers—from psychotherapists and counselors to social workers and peer supporters [2, 3]. While there is increasing interest in exploring direct-to-patient AI systems for some circumstances, and some promising results [4], we expect the demand for human therapists and mental health support to continue to far exceed demand. The limited supply of effective therapy providers is due—at least in part—to the reliance on resource-intensive methods to train helping skills [5] and evidence-based interventions [6, 7]. Traditional methods require access to trainers who can simulate a client interaction [8, 9] and provide expert supervision [10, 11], creating significant barriers to growing high-quality training [12, 13].
Fortunately, AI systems have been increasingly applied to counselor training as a potential solution to these scaling challenges. Recent advances in large language models (LLMs) have enabled the development of LLM-based systems for simulating patients seeking mental health support [14, 15], offering rich opportunities for practice. Simultaneously, AI feedback systems have progressed in automatically assessing counselor behaviors such as empathy, reflections, and active listening [16, 17, 18, 19] and generating suggested responses [20, 21, 22] and explanations [23, 22]. These AI feedback systems target skills from client-centered approaches [24, 25]—empathy, reflections, questions, and active listening—which are relevant to all mental health providers. Meta-analyses have shown these skills strengthen common factors like therapeutic alliance, which are powerful predictors of therapy outcomes, regardless of the eventual evidence-based intervention or therapy modality implemented [26, 27, 28]. In addition to technical advances, prior work evaluating AI feedback systems with novice counselors have tested their effectiveness when used as real-time co-pilots rather than as training systems [29, 21], or evaluated pre-LLM training systems which had limited practice realism and simpler, non-generative feedback mechanisms [30]. Therefore, we have limited evidence explaining how modern, LLM-based training systems—which offer practice simulations resembling challenging real-world behaviors and generative feedback with suggested alternatives and rationales—promote counselor skill development [31, 32]–a holistic concept that includes appropriate use of skill behaviors, accurate self-awareness of one’s capabilities and limitations [33], and an ability to reflect on and learn from experiential practice [34].
To address this gap, we develop CARE, an LLM-based training system for novice counselors, and conduct a randomized experiment to investigate how two different modes of simulated training impact counselor skill development. CARE enables (1) realistic practice with LLM-simulated patients prompted by expert counselors to resemble challenging behaviors [15], and (2) structured feedback from a fine-tuned LLM that identifies strengths and areas for improvement across core counseling skills (e.g., empathy, reflections, questions, suggestions), while also providing explanatory rationale and alternative responses [22]. Importantly, CARE’s feedback evaluates not only when a counseling skill is used but also how well it is used, distinguishing between strong implementations (e.g., reflections that accurately capture client emotions) and implementations needing improvement (e.g., reflections that miss core client concerns). These feedback evaluations are based on established counseling frameworks [5, 35] and expert-annotated examples.
Figure 1: Our between-subjects study randomizes participants to either practice with AI Patients alone (P) or practice with AI patients and receive AI feedback (P+F). We holistically evaluate counselor skill development from three perspectives: automatic assessments of behaviors of skills used via LLM classifiers; self-efficacy and its calibration with actual performance; and qualitative self-reflections after the training intervention chat and post-intervention chat.
We conducted a 75-minute online lab study with novice counselors (N=94) to empirically evaluate how different LLM-simulated practice modes in CARE impact skill development. Participants were randomly assigned to one of two conditions: (1) Group P: Practice with LLM-simulated patients without AI feedback, or (2) Group P+F: Practice with LLM-simulated patients plus AI feedback (see Fig. 1). To comprehensively study learning outcomes, we measured changes from pre-intervention to post-intervention across multiple dimensions: actual behavioral performance (calculated via automatic analysis of counseling transcripts), participants’ self-efficacy (perceived confidence in counseling skills), and metacognitive awareness (captured through guided self-reflection prompts). Our study design specifically addressed: What changes occur after practice with an AI-simulated patient alone? How do these outcomes differ when participants also receive structured AI feedback?
Our findings reveal that practice with feedback significantly improves empathetic listening skills, while practice alone shows minimal improvement in these core competencies. Notably, both groups failed to improve the accuracy of participants’ self-efficacy assessments, highlighting the challenges involved in helping learners develop a well-calibrated understanding of their own performance. Through analysis of participants’ self-reflections, we found the P+F group internalized the importance of empathetic and active listening, but P participants continued to overly focus on solutions, albeit with increased information-gathering. These insights suggest that future LLM-based training systems must integrate feedback mechanisms to cultivate the empathetic stance fundamental to effective counseling.
Stages Skill Category Description
Questions Questions seek information from the client and can be open (inviting elaboration) or closed (requesting specific answers). They include both direct questions and indirect prompts (e.g., ”Tell me about…”).
Reflections Reflections capture and return to clients something they have communicated, either explicitly or implicitly. They typically mirror back content from the client’s preceding statement, but can also reference earlier parts of the conversation.
Empathy Empathy can be shown through emotional warmth, interpretation of the client’s experience (e.g., paraphrasing, making conjectures, or sharing relatable experiences), or exploration of the client’s feelings and perspectives.
Suggestions Suggestions offer possible actions, perspectives, or solutions in a respectful and autonomy-supportive manner. They may involve information-sharing or proposing alternative viewpoints.
Session Management Session management includes organizing the session, transitioning between topics, and summarizing key points. It provides structure and helps maintain therapeutic focus.
Table 1: Our analysis of skill development has a theoretical-foundation in Hill’s Helping Skills model [5]. We select a skill subset relevant for beginning counselors at the undergraduate and first-year graduate level [36]. These include microskills during the exploration and action stages; and macro skills that are applicable throughout the session. Hill’s insight stage, of which self-disclosure was the only relevant skill for basic counseling, was excluded from our primary analyses due to its infrequent occurrence in our data.
1 Results
This section presents findings from our investigation of how LLM-based practice affects counselors’ skills development. To understand the impact of simulated practice alone (P) and practice with feedback (P+F), we integrate evidence from three complementary data sources.
Behaviors of Skills Used: Training systems should help improve the use of skills that align with evidence-based counseling practice. Thus, we analyzed the skills used in transcripts and how they change between the pre-intervention and post-intervention assessments using high-accuracy LLM classification models. We assessed behaviors of skills used for the exploration stage (strong uses in empathy, reflections, questions) and action stage (suggestions needing improvement) of Hill’s Helping Skills framework [5]; see Table 1 for definitions.
Self-Efficacy Measures: It is informative to understand participants’ confidence and their awareness of their own strengths and limitations. We measured counselor self-efficacy using the Counselor Activity Self-Efficacy Scale [37], which covers confidence for exploration skills, action skills, and session management skills (Table 1). Our analysis tests whether the lowest performing individuals overestimate their abilities (aka the Dunning-Kruger effect [38]) and whether LLM-training interventions improve the accuracy of their self-assessments and raw self-efficacy scores.
Qualitative Data on Learner Takeaways: LLM-simulated training provides opportunities for experiential learning opportunities [39] whereby reflection on action [34] can support updates to one’s knowledge. We also asked participants’ open-ended self-reflections following the practice intervention and post-intervention assessment, and qualitatively coded their responses.
We integrate these three data sources to provide a holistic view of how AI-based practice—with and without AI feedback—influences changes in counselors’ use of skills, self-efficacy, and key lessons learned.
1.1 Changes in Counseling Skill Behaviors
Our primary analysis uses LLM-based behavioral assessments of transcripts to evaluate whether skill usage changes from pre-intervention to post-intervention for both the practice and practice-and-feedback groups. In a secondary analysis, we investigate how exposure to good examples of skills within the AI feedback may explain any changes we observe.
Change after P Change after PF Differences in
change after P vs. P+F
[…]
Empathy ($\uparrow$) -9.6 0.001 -0.52 5.4 0.117 0.23 15 0.001 0.72
Reflections ($\uparrow$) 1.4 0.391 0.18 3.7 0.034 0.32 2.3 0.323 0.2
Questions ($\uparrow$) 2.4 0.421 0.12 6.6 0.018 0.36 4.1 0.296 0.22
Suggestions ($\downarrow$) -5.9 0.010 -0.39 -5.5 0.057 -0.28 0.4 0.910 0.02
Figure 2: Changes in counseling behaviors following AI patient simulations alone (P) versus AI patient simulations with AI feedback (P+F). The plot displays bootstrapped means for pre-intervention and post-intervention interactions. The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes). Notably, P group experiences a significant drop in strong uses of Empathy (-9.6% change, $d=-0.52$), whereas P+F group’s use of Empathy is maintained and trends towards improvement. The P+F group also experiences noticeable improvements in Reflections (+3.7% change, $d=0.32$) and Questions (6.59% change, $d=0.36$)
We find that practice alone is not enough; feedback during practice is necessary to promote desirable counseling behaviors in empathetic and active listening. AI feedback during practice (P+F) led to improvements in Reflections (+3.6% change, $p=0.034$) and Questions (+6.59% change, $p=0.018$), and trended towards improvement in Validation (+3.6% change, $p=0.083$), Suggestions (-5.45% change, $p=0.057$) and Empathy (+5.37% change, $p=0.117$). In contrast, practice alone (P) led to only improvements in Suggestions (-5.85% change, $p=0.011$), but no improvements in Reflections, Questions, and Validation, and significantly worse expression of Empathy (-9.6% change, $p<0.001$). Pairwise comparisons between the two conditions were substantial and significant for Empathy (15% relative difference, Cohen’s $d=0.72,p<0.001$) and non-significant for others.
To better understand AI feedback’s role, we further analyzed whether the level of exposure to specific feedback during practice affected these changes. Our analysis quantifies the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length. We hypothesized that greater exposure to GAP for a skill would lead to stronger use of that skill. To test this, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$).
[…]
Figure 3: Counselor Self Efficacy (perceived ability to use skills) for participants grouped by behaviors of skills used (actual performance). Notes: Gaps depict miscalibration between actual and self-assessed percentile of performance for quartile groups with bootstrapped 95% CIs. We only visualize for data collected in the post-intervention.
1.2 Self-Efficacy and its Miscalibration with Behaviors
Our investigation of self-efficacy involved three types of analyses. First, we examine the calibration of self-efficacy ratings with actual performance and evaluate whether P or P+F interventions improve this. Second, we tested for significant changes in raw self-efficacy scores after practice or practice-and-feedback.
Our analysis comparing self-efficacy ratings with actual performance across skill quartiles finds support for the Dunning–Kruger effects more substantially for action skills and to a lesser degree for exploration skills. The interaction between measure and quartile was significant in four out of six ANOVAs for action skills, while only one out of six was significant for exploration skills (Table 4). Pairwise comparisons also showed a pattern indicative of a Dunning–Kruger effect (see Table 5, Table 6 and Fig. 3): People in the lowest quartile overestimated themselves the most. Those in the highest quartile—and to a lesser degree also those in the second-to-highest quartile—tended to underestimate themselves.
Change after P Change after PF Differences in
change after P vs. PF
[…]
Exploration Skills 0.36 0.004 0.44 0.13 0.338 0.14 -0.21 0.238 -0.25
Action Skills 0.27 0.166 0.21 0.33 0.026 0.34 0.04 0.884 0.03
Session Management 0.35 0.021 0.35 0.36 0.011 0.39 0.01 0.955 0.01
Figure 4: Changes in raw-scores of self-efficacy following AI patient simulations alone (P) versus AI patient simulations with AI feedback (PF). The plot displays bootstrapped means for pre-intervention and post-intervention. The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes).
Participants’ Ability to Self-Assess Their Skill Level Remained Mixed After LLM Practice. For the practice only (P) group, the mean discrepancy in exploration skills changes from 11.6 percentile underconfidence to 5.7 percentile overconfidence, a significant shift ($p<0.001,d=0.58$). Besides this, we found no other significant changes in calibration. No significant differences was found among the P group for discrepancy in action skills ($p=0.227,d=0.18$). The P+F group showed no significant calibration changes for exploration skills ($p=0.479,d=-0.10$) or for action skills ($p=0.393,d=0.13$). Finally, between-group differences were not significant for discrepancy in exploration skills ($p=0.191,d=0.27$) or action skills ($p=0.743,d=0.07$).
We do find modest overall increases in self-efficacy after P and P+F interventions, with different patterns of improvement across skills (Fig. 4). For the P group, confidence in exploration skills showed a significant increase ($0.36$ points on an 8 point scale, $d=0.44,p=0.004$). Confidence in session management skills showed a significant increase for the P+F group ($0.36$ points, $d=0.39,p=0.011$). While session management skills for the P group also trended towards improvement, it was not significant after correcting for multiple hypothesis testing ($0.35$ points, $d=0.35,p=0.021$). Similarly, while confidence in action skills for the P+F group also increased, this result was not significant after correction of multiple hypothesis tests ($0.33$ points, $d=0.34,p=0.026$). Finally, we found no significant differences between participants who received AI feedback (P+F) versus those who did not (P) (across the three self-efficacy subscales, $d=-0.25,0.03,0.01$, $p=0.238,0.884,0.955$).
1.3 Qualitative Self-Assessments after AI practice with and without AI feedback
To understand participants’ learning process during CARE practice, we analyzed qualitative self-reflections from two time points: immediately after the training intervention chat, where participants responded to “What would you do differently as a therapist?” and after the post-intervention chat, where they reflected on “What did you do well as a therapist?”. We examined how initial intentions translated into reported strengths across the P+F and P groups. Two high-level themes emerged in how participants approached skill improvement:
(1) The P+F group expressed greater intentions and successes in improving their use of empathy and listening skills. P+F participants reported effectively using empathy (27%), validation (27%), and open-ended questions (52%). They emphasized the value of listening skills, such as reflective responses to signal understanding: “I should rephrase what they say so they know I’m understanding them” (P51). They also recognized that counseling should support client exploration of thoughts and emotions, rather than provide direct solutions. One participant reflected on this shift: “I asked them to expand on their feelings, rather than guiding them to my idea” (P39). AI feedback encouraged this shift toward providing emotional support and fostering client autonomy, helping counselors adopt a more empathetic and client-centered approach.
(2) The P participants remained solution-oriented but changed their approach to first gather information. While P+F participants intended to give fewer suggestions, many P participants continued to view suggestions as a central skill. In the post-intervention, 48% of P participants reported using suggestions successfully, compared to only 14% of P+F participants. Many P participants justified their continued use of suggestions by citing a desire to provide tangible, actionable help, for example, “Maybe because I’m untrained and solution oriented. I do not want to leave them with nothing, and nowhere to go” (P40). Some reflected on modifying how they delivered suggestions, emphasizing strategies like gathering more information to tailor advice or providing more concrete guidance. However, in the absence of feedback, most remained fixed in their approach, with some even reporting efforts to rephrase the same solution repeatedly to persuade the client.
2 Discussion
This study investigated the impact of practicing with an LLM-simulated patient either with or without receiving LLM-generated feedback on counselor skills development, resulting in three main findings. First, our behavioral assessments showed that practice with feedback improves empathetic listening skills, while practice alone shows minimal improvement, as evidenced by decreased use of empathy. Second, our qualitative analysis of self-reflections revealed distinct skill development strategies, with feedback recipients more frequently reporting the adoption of client-centered approaches focused on showing empathy and exploring patients’ thoughts and feelings, while practice-only participants gravitated toward solution-oriented approaches focusing on gathering more information and providing suggestions. Both these findings highlight that the development of counseling skills requires not only practice opportunities but also structured feedback that guides learners toward empathetic, client-centered approaches. Finally, participants demonstrated poor calibration between their perceived abilities and actual performance, as evidenced by overestimates of self-efficacy for the lowest quartile performers. This underscores how self-efficacy measures may not reliably indicate skill development. Each of these findings merits further discussion.
Our findings demonstrate that teaching counselors to implement client-centered approaches requires more than just simulated practice opportunities—it requires targeted feedback that guides novices away from their natural solution-oriented tendencies. The fully-featured version of CARE–combining LLM-simulated patient practice and LLM-based feedback–helped participants improve their use of empathetic and active listening skills, with notable improvements in questions ($d=0.36$) and reflections ($d=0.32$). In comparison, practice with an AI patient alone only led to fewer inappropriate suggestions ($d=-0.39$), but no improvements in reflections or questions, and significantly worse uses of empathy (-9.6% change, $d=-0.52$; 15% relative difference to P+F, $d=0.72$). These effect sizes are comparable to those found in studies of human supervision during standardized roleplays, where Maaß et al. (2025) [41] reported observer-rated skill improvements with effect sizes ranging from $d=0.29-0.49$. Since LLM-simulated practice and feedback are not bottlenecked by the resource constraints of human trainers, AI training systems like CARE show promise in scaling access to effective counseling training.
While the combination of feedback and simulated practice was key to CARE’s success, we also recognize that CARE represents just one approach to providing expert-aligned LLM feedback to learners. CARE allows trainees to request feedback on responses they have already sent. This approach allows trainees to experience productive learning discomfort by trying and making mistakes in simulation, while still offering feedback upon request to refine future responses. This design choice parallels elements of both live human supervision [42] and delayed supervision after a session [41], creating a hybrid approach that may offer flexibility based on trainee preferences. Given the promise of feedback combined with practice, the AI for psychotherapy field should evaluate different AI-feedback designs against each other. Several other possible strategies for giving feedback include just-in-time skill suggestions [21, 43], counterfactual simulations [44], and global session feedback. While conducting randomized trials in medical education research can be tricky [45], randomized trials delivering different training interventions to large participant-pools for longer durations [46] could provide essential data to determine optimal approaches for different learning objectives and learner characteristics.
Our study revealed that participants’ self-efficacy ratings were poorly calibrated with their actual performance, especially among lower performers. This finding, consistent with prior research, suggests that self-assessment accuracy alone may not be a reliable indicator of counselor competence or development. Recent reviews indicate that efforts to improve self-assessment calibration have limited impact on learning or performance outcomes [47]. Instead, it is valuable to objectively assess specific standards of performance and skill use [42, 48] and design interventions that can help low performers improve on those metrics while maintaining a positive morale for continued practice. As AI-based training tools evolve, integrating objective performance measures and structured self-reflection, rather than relying solely on self-assessment, offers a more robust approach to supporting counselor development.
Several limitations of this study warrant acknowledgment. First, our assessment focused primarily on immediate skill acquisition rather than long-term retention or transfer to actual clinical settings. Second, while our automated behavioral analysis focused on microskill usage, it could not capture observable session-level characteristics. While traditional studies have employed human observers to provide such ratings, recent AI research has shown the validity of using finetuned LLMs in certain contexts to approximate session-level measures, such as working alliance [49], which might enable scalable behavioral analyses of more skill development constructs. Third, we recruited from a sample of bachelor-level counselors through the online Prolific platform, which may differ from in-person or classroom settings. For example, existing training environments for counselors will typically offer didactic materials (e.g., lectures and textbooks) and non-AI training opportunities (e.g., peer-to-peer roleplay and feedback). Thus, future work should evaluate the utility and acceptability of AI-based training tools within existing educational and training pathways. Finally, the efficacy of LLM-based practice and feedback training was only shown for client-centered microskills. Future work is required to understand whether these results generalize to other therapy modalities (e.g., psychodynamic, CBT/ACT), where training interventions may work to shift natural tendencies in similar or subtly different ways.
3 Conclusion
Despite increasing interest in using LLMs in mental health, to our knowledge, this is the first study to conduct a large-scale evaluation ($N=94$) of an LLM-based training system for developing core skills in novice counselors. Our findings show that while simulated practice (P) proved insufficient—and in the case of empathy, potentially harmful—practice with AI-generated feedback (P+F) supported measurable improvements in key counseling skills of demonstrating empathy, delivering reflective responses, and engaging in client-centered inquiry. By combining realistic patient simulations with expert-aligned, skill-specific feedback, the CARE AI Training system can help novices to master skills involved in client-centered therapy, offering a promising avenue for scalable, evidence-aligned training in mental health care.
4 Methods
4.1 Overview of CARE Training Platform
Figure 5: CARE’s practice and feedback model visualized in a diagram and web screenshot. In CARE, counselors practice with an LLM-simulated patient and receive feedback on each of their responses. The feedback model labels whether a response has strengths or constructive feedback areas. Responses with constructive feedback explain what the goal should be at this point in the conversation; what a helper could improve to better align with this goal; and how they could respond differently via an alternative response.
We developed CARE as a web platform for novice counselors to train in text-based counseling skills enabled by LLMs. Trainees practice simulated conversations with AI patients and review feedback on their responses generated by an AI mentor. CARE builds on top of successes from previous research in co-designing with mental health experts to improve the realism of LLM-simulated patients [50, 43, 14, 15] and finetuning domain-specific LLMs with therapeutic knowledge capable of generating feedback and alternative responses for text-based peer counseling conversations [51, 22, 52, 53, 54].
CARE allows novice counselors to develop their counseling skills in a text-based format by practicing with AI-simulated patients and receiving feedback on their responses (see Fig. 5).
Consider Aki, a novice peer counselor who wants to use CARE to experience hands-on training using counseling skills that they have recently read about. In CARE, Aki can practice with an AI patient of their choice from a library of patient scenarios. Aki starts a practice chat with an AI patient with the background of a 35-year-old male veteran seeking to reconnect with his children but facing legal barriers and parental gatekeeping.
Each practice scenario provides limited background information about the AI patient and their presenting problem (e.g., ”Young adult with family issues: low mood and self-esteem”). This intentional limitation requires counselors to simultaneously learn more about the patient’s situation while demonstrating empathy and support.
Aki starts the conversation by greeting the AI patient. The AI patient shares their frustration with feeling everyone is against him, and hopes to find ways get his kids back and overcome the difficulty his judgmental parents. Aki composes a reply, and the conversation continues in this turn-by-turn manner.
Users can view feedback on their therapeutic responses at any point during or after the session, offering flexibility to either review intermittently throughout the practice or comprehensively after completion.
4.2 Implementation details for CARE Training Platform
CARE was built as a web application with a Python Flask back-end and React Javascript front-end which online participants could access and interact with. LLM-simulated patients in CARE were powered by the GPT-4o API with a custom prompt defining the patient’s background scenario and behavioral principles [15]. The prompts for the three AI patients used for the pre-intervention assessment, intervention, and post-intervention are provided in the Supplementary Materials A.1. The LLM feedback system was powered by a Llama-2 13B parameter model finetuned on an expert-annotated feedback dataset [22].
4.3 Study Procedures
Between August and October 2024, we conducted an online lab study with (N=94) novice counselors. Given our aim to train novices lacking formal counseling training, we recruited participants who expressed an interest in practicing helping skills yet had minimal prior education or practical experience. Specifically, we used the Prolific platform to screen for participants who (1) lacked extensive formal professional or graduate training in clinical psychology, counseling, social work, or nursing (excluding those who were pursuing or had completed master’s or doctorate degrees); and (2) had less than one year of counseling-related experience (e.g., peer support or crisis counseling). In our final experiment, we recruited N=94 participants. Additionally, during the pilot phase, we recruited another N=14 participants who were allowed broader educational and hands-on experiences for protocol refinement, but we report qualitative and quantitative results exclusively from the final set of 94 participants.
Each participant completed a 75-90 minute video-call session following the structure outlined in Fig. 1. After receiving a static 5-minute tutorial on core counseling skills (see Fig. 6), all participants completed a 10-minute pre-intervention assessment chat with the first AI patient. For the 20-minute training intervention, participants were randomly assigned to either practice with an AI patient only or practice with the option to receive feedback on all of their sent responses. Finally, all participants completed a 10-minute post-intervention assessment chat with a third AI patient. Participants answered self-efficacy items after the pre-intervention and post-intervention chats, and provided qualitative self-reflections after the training intervention chat and post-intervention chat. All participants were paid \$15/hour. To protect participant privacy, we instructed them to join the video-call using their ProlificID as the display name, turn off their video-camera, and share their screen. Participants consented for the video call to be recorded for our own analysis and their CARE usage data to be used and shared for the purposes of this and future research after de-anonymization. This study protocol was approved by our institution’s IRB.
4.4 Assessing Changes in Skill Use through LLM Classifiers
We investigate whether counselors employ higher-quality counseling behaviors after LLM-based training and between training conditions by leveraging NLP methods. This analysis is motivated by the need to quantify changes in counseling skill use at scale across multiple participant sessions. Our approach follows four key steps: (1) fine-tuning and validating LLM-based classifiers to identify skill behaviors, (2) selecting a final set of classifiers based on performance metrics and theoretical priority, (3) testing changes in skill use across time and between intervention groups, and (4) analyzing the relationship between exposure to good alternative patterns in AI feedback and subsequent skill use.
4.4.1 Fine-tuning and Validating LLM-based Classifiers
We developed LLM-based binary classifiers that allows us to label the skill use within a transcript. For example, one of our fine-tuned classifier could determine which utterances in a transcript had strong uses of Questions during that stage of the transcript. To finetune and evaluate these classifiers, we transformed a previously published expert-annotated, feedback dataset that had strengths and areas of improvements for 8 skills [22] into a 16-class binary classification format (8 skills × 2 categories each—strong uses and areas needing improvement) [^1] The binary classification feedback dataset can be accessed at https://huggingface.co/datasets/youralien/feedback_qesconv_16wayclassification. Additionally, we used a subset of transcripts from this randomized study which were further annotated by counseling domain-experts.
Three experts were recruited with relevant backgrounds including practicing clinical psychologist, licensed marriage family therapist, former director and supervisor of a crisis agency. Each of the experts annotated 10 participants’ study transcripts (5 from the practice-only group; 5 from the practice-and-feedback group), totaling 370 counselor utterances. Two rounds of annotation occurred: after collecting a first annotation pass and identifying data points with disagreements, we showed each of the experts the others’ annotations and had them re-annotate and provide rationales for their decision. We display pairwise agreement results averaged across all pairs in Table 2. Note that while we originally explored other annotation agreement metrics such as Cohen’s kappa, the severe class imbalance of our annotations made these metrics less relevant in our case. The gold-standard CARE 10% expert-annotated sample consists of labels that result from a majority vote across these three experts [^2] This expert-annotated data sample can be found at https://huggingface.co/datasets/youralien/CARE_10percent_16wayclassification.
We finetuned RoBERTa-large binary classifiers using a 95% split of the transformed FeedbackQESConv dataset, and did hyperparameter tuning on 5% of FeedbackESConv and the CARE 10% sample. The performance of the our LLM-based classifiers are shown in Table 2. The highest performing classifiers ($F1>0.5$) became candidates for our automatic behavioral assessments, which we further down-selected as described below.
Skill Strengths Areas to Improve
Annotator Classifier Annotator Classifier
Agreement Performance Agreement Performance
% acc. f1 % acc. f1
Empathy 0.793 0.813 0.741 0.809 0.859 0.389
Reflections 0.863 0.900 0.562 0.944 0.903 0.312
Questions 0.732 0.784 0.775 0.852 0.842 0.394
Suggestions 0.919 0.955 0.507 0.946 0.941 0.681
Validation 0.726 0.852 0.556 0.919 0.893 0.265
Self-disclosure 0.982 0.920 0.326 0.969 0.986 0.849
Session Management 0.968 0.941
Professionalism 0.905 0.969
Table 2: Left columns show pairwise annotator-agreement averaged across 3 domain-experts for the CARE 10% sample (n=370). Right columns show performance of the best RoBERTa-large classification models after hyperparameter tuning on our validation dataset, CARE 10% sample (n=370) + FeedbackQESConv 5% sample (n=409). Session Management and Professionalism were excluded from finetuning due to infrequent occurrence.
4.4.2 Down-selecting a Final Set of Classifiers
From the initial set of 16 binary classifiers, we applied both methodological and theoretical criteria to select our final set for analysis. First, we established a minimum performance threshold of $F1>0.5$ to ensure reliable classification. This criterion yielded seven candidate classifiers: strong uses of Empathy, Reflections, Questions, and Validation, as well as both strong uses and areas needing improvement for Suggestions.
To maintain statistical power while controlling for multiple comparisons, we further narrowed our focus to four key classifiers: strong uses of Empathy, Reflections, and Questions, plus areas needing improvement for Suggestions. This selection was guided by three considerations: (1) the need to limit the number of statistical comparisons to avoid diluting significance across too many tests, (2) focusing on classifiers with the strongest performance metrics, and (3) selecting skills that represent core competencies in client-centered frameworks and are frequently used in counseling sessions.
Self-disclosure was excluded from our analysis due to its infrequency in our dataset, while Validation, though conceptually related to Empathy and mentioned frequently in qualitative data, showed more limited classifier performance and was therefore reserved for secondary analyses.
4.4.3 Testing Changes across Time and between Groups
Using our selected classifiers, we examined how skill usage changes from the pre-intervention chat to the post-intervention chat. For each session transcript, we computed an aggregate score defined as the proportion of utterances with strong uses of a skill (e.g., $b=N_{strengths}/N_{total}$) or with areas needing improvement (e.g., $b=N_{improvement}/N_{total}$).
To statistically test for changes between the pre-intervention and post-intervention chats ($b_{0},b_{1}$), we used paired t-tests and computed Cohen’s $d$ effect sizes. To test for significant differences between P and P+F groups ($b_{1}^{P}-b_{0}^{P}$ vs. $b_{1}^{PF}-b_{0}^{PF}$), we used non-paired t-tests.
In total, we conducted 12 planned t-tests for analyzing changes in behaviors: three skills (Empathy, Reflections, Questions) with strong uses and one skill (Suggestions) with areas needing improvement, each analyzed for both within-group changes and between-group differences in those changes.
4.4.4 Exposure to Good Alternatives in AI Feedback
To better understand AI feedback’s role in skill development, we analyzed whether the level of exposure to specific feedback during practice affected post-intervention performance. We quantified the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length.
To test whether greater exposure to GAP for a skill led to stronger use of that skill, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$). This model allowed us to isolate the effect of feedback exposure while accounting for baseline performance differences.
4.5 Analyzing Confidence and the Calibration of Self-Assessments using Counselor Self-Efficacy
Counselor self-efficacy represents a critical dimension in counselor development, defined as one’s “beliefs or judgments about their ability to effectively counsel a client” [55]. While high self-efficacy can promote persistence and resilience when facing challenges [56], miscalibrated self-assessments—particularly overconfidence relative to actual ability—can lead to suboptimal or potentially harmful counseling behaviors [57, 58, 59].
This calibration challenge has prompted research into potential remedies. Prior studies suggest that increased practice experience may help individuals better align their self-assessments with actual performance, though results remain inconsistent [60]. Similarly, supervisor feedback has been proposed as a mechanism to address self-assessment miscalibration [61], although evidence from other educational contexts shows mixed effectiveness [62]. Our methodology specifically examines whether practice alone or practice with structured feedback differentially impacts both raw self-efficacy and its calibration with actual performance.
4.5.1 Self-Efficacy Measurement Instruments
To measure counselor self-efficacy, we employed the Counselor Activity Self-Efficacy Scale (CASES) [37], specifically utilizing a revised subset of items targeting basic counseling skills (CASES-R) [63, 36]. The CASES-R established a three-factor structure to assess counselors’ confidence in performing key therapeutic functions: Exploration and Insight Skills (EIS), Action Skills (AS), and Session Management Skills (SM).
Participants completed the CASES-R immediately following both pre-intervention and post-intervention AI patient interactions. All items were administered using an 8-point Likert scale (0 = no confidence, 7 = complete confidence). During factor analysis, we discovered that among the five original EIS skills, self-disclosure did not load on the same factor as the other items. Consequently, we consolidated the EIS dimension to include only four Exploration Skills: Reflections of Feelings, Restatements, Open Questions, and Listening.
The final instrument comprised 12 items across three factors: (1) exploration skills (e.g., restatements, reflecting feelings, open questions, listening); (2) actions skills (e.g., providing suggestions, knowing which actions to take); and (3) session management skills (e.g., keeping sessions on track). Reliability analysis demonstrated good to excellent internal consistency across all factors, with Cronbach’s $\alpha$ values of 0.784, 0.803, and 0.905 for ES, AS, and SM, respectively.
4.5.2 Dunning-Kruger Effect Analysis
Our primary analysis investigated potential mis-calibration between participants’ self-assessments and their actual counseling performance, specifically examining whether data exhibited patterns consistent with the Dunning-Kruger effect. This phenomenon [38] suggests that individuals with lower skill levels tend to overestimate their abilities, while highly skilled individuals may slightly underestimate their competence. We focused this analysis on Exploration Skills (ES) and Action Skills (AS), as these dimensions had straightforward mappings between CASES items and our NLP behavioral classifiers (Table 3).
To test for the presence of the Dunning-Kruger effect, we follow the classic analysis method that splits the data into quartiles based on performance and conducts a two-way analysis of variances for self-assessments and actual performance and across the quartiles; and finally verifies via post-hoc tests that the bottom performers have the biggest overestimation of their abilities [38]. To standardize the comparison between self-efficacy and performance, we transform each measure into a percentile rank (0 - 99) computed across all data collected for the pre-intervention and post-intervention chats ($b_{0},b_{1}$; $s_{0},s_{1}$).
4.5.3 Calibration Change Analysis
To evaluate whether our interventions improved self-efficacy calibration, we computed discrepancy scores by subtracting standardized performance scores from standardized self-efficacy scores for each participant at both assessment timepoints. These discrepancy metrics provided a direct measure of calibration, with positive values indicating overconfidence and negative values indicating under-confidence.
We then examined changes in these discrepancy scores from pre- to post-intervention for both intervention groups. This analysis allowed us to determine whether either intervention significantly improved the alignment between participants’ self-perceptions and their actual counseling abilities.
4.5.4 Changes in Raw Self-Efficacy
Beyond calibration, we also investigated whether the interventions affected participants’ absolute levels of self-efficacy across the three measured dimensions (ES, AS, and SM). We conducted repeated-measures analyses to identify: (1) significant pre-post changes in raw self-efficacy scores following practice alone (P intervention); (2) significant pre-post changes in raw self-efficacy scores following practice with structured feedback (P+F intervention); and (3) differential patterns of change between the P and P+F groups, indicating potential intervention-specific effects on self-efficacy development.
4.6 Interpreting Counselor Learning through Self-Reflections
For qualitative analysis, three authors coded interview transcripts using deductive thematic analysis [64] to understand participants’ experiences with CARE. To manage the 94 transcripts efficiently, we used timestamped session notes to identify relevant observations, referring back to the video recordings for verbatim quotes when necessary. After two rounds of codebook refinement, three co-authors coded these data points to inform our qualitative findings.
We analyzed two types of qualitative self-assessment questions. A set of low-level responses (e.g., “What did you do well?” and “What would you do differently?”) were deductively coded using the 8 counseling skill categories defined in the CARE feedback framework [22], allowing us to assess alignment between self-reported reflections and the system’s feedback focus. A second set of high-level responses was coded inductively, identifying emergent patterns in participants’ reflections and lessons learned.
We structured the final thematic findings by comparing how participants in the P group and the P+F group differed in their self-reported skill use, intentions for improvement, and reflections on therapeutic mindset. Specifically, we separated responses made by participants in the P and P+F groups, and then analyzed differences in the frequency and framing of each coded skill (e.g., empathy, validation, suggestion-giving).
5 Contributors
Ryan and Raj led the technical development efforts of CARE, the LLM-simulated training system. Ryan, Emma, and Diyi worked together to design the randomized experiment. Ryan, Ifdita, Juan Pablo, and Raj conducted the 90+ online video call sessions. Ryan was responsible for the automatic behavioral assessment of counseling skills, as well as the self-efficacy calibration analysis, while Emma and Diyi advised on analysis methods and interpretations of these quantitative results. Ifdita, Juan Pablo, and Ryan conducted the thematic analysis of participants qualitative, self-reflection data. All authors helped with drafting and editing the Article manuscript and figures, with a huge thanks for Raj, Diyi, and Emma for completing full editing passes over the Article.
6 Acknowledgments
This research was made possible with funding support from a Stanford Impact Labs Grant, Stanford HAI Seed Grant, Stanford Psychotherapy and Behavioral Sciences Department Innovator Award, and Stanford CURIS paid-internship program for undergraduate researchers.
The authors have a large community of researchers to thank for their help during all stages of this project. The CARE system would have not been possible without the significant code contributions from student researchers in SALT Lab, including Meijin Li, Cheng Chang, Ananjan Nandi, Alicja Chaszczewicz, and Alan Zhang. In addition, CARE’s design benefited from feedback from Bruce Arnow and others from the Stanford-PAU PsyD consortium. We thank Pannisy Zhao for assistance while conducting user studies. Statistical analyses and presentation of results were much improved due to help from Robert Kraut and Akhila Kovvuri. Thank you Yanzhe Zhang and Rose Wang for detailed discussions on finetuning the LLM classifiers. Thank you to the following individuals for providing feedback on early drafts and presentations of this work, with special thanks to Yutong Zhang, Will Held, Ella Li, Michael Ryan, Dora Zhao, Caleb Ziems, Matthew Jöerke, Joy He-Yueya, Omar Shaikh, and Kapil Garg.
Appendix A Supplementary Materials
A.1 Prompts for CARE’s LLM-simulated Patients
The following patients were used during the randomized online lab study. Participants interacted with the same ordering of three AI patients in the pre-intervention assessment chat, intervention chat, and post-intervention chat.
These three patients were previously tested with domain-expert judges [15] and were rated an average of 6 out of 7 for realism. They were chosen to balance (1) diversity across different ages, genders, and presenting counseling concern; and (2) consistency in challenging behaviors. In particular, each patient’s prompt defined similar behavioral principles to either not suggest solutions on one’s own, or to show a degree of skepticism or reluctance to a therapist’s advice or solutions. Thus, these AI patients were designed to challenge the novice counselors in similar ways, despite the diversity in scenarios. We highlight in blue these principles that resemble this resistance to arriving at or accepting solutions.
A.1.1 AI Patient for Pre-intervention Assessment Chat
**Name and Bio**:
35-year-old American male: Feeling Alone After a Holiday
**Scenario:**
You are a 35-year-old American male. You are feeling abandoned and alone after the holidays. Everyone had been with family but you are not talking to your parents. You feel the injustice of being abandoned and have no interest in an olive branch to work on things.
**Principles to adhere to:**
1. Keep your responses short and to the point
2. You limit your replies to 1 - 3 sentences.
3. Feel free to make up believable stories about your past to answer any questions
4. Do not repeat sentences or the same emotion words.
5. When presented with suggestions, show a degree of skepticism or reluctance to accept the advice immediately. This can be done by questioning the feasibility of the suggestion or by expressing uncertainty about whether it’s the right solution for you.
6. When expressing doubts or fears, avoid jumping to solutions. Instead, articulate the concerns and allow the conversation to explore these feelings more deeply@
7. Dont be so self-aware or good at recognizing your own problems@
8. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
A.1.2 AI Patient for Intervention Chat
**Name and Bio**:
35-year old Male Veteran: Substance use and legal issues
**Scenario:**
The member is a 35-year-old male, cisgender, heterosexual
veteran who has recently presented to treatment to address
his substance use issues and legal issues. He is court-
mandated to therapy. He had a severe psychotic break in
the context of marijuana and psilocybin about 6 months ago,
where he experienced paranoid delusions and hallucinations.
He is now stable and is not experiencing any psychotic
symptoms. In therapy, he is hoping to work on his estranged
relationship with his parents, who are currently caring for
the members two young children. The member is adamantly
focused on being reunited with his parents. In therapy, the
member is unable to accept that he had a recent episode of
psychosis and is very resistant to anything that resembles
criticism. He does not view himself as having any issue and
believes that all of the problems in his life are because
of other people. His demeanor is hostile and somewhat
aggressive, and he is quick to shut down any conversation
that might identify his own areas of development. He
struggles to feel emotions beyond anger and frustration.
**Principles to adhere to:**
1. Keep your responses short and to the point
2. You limit your replies to 1 - 3 sentences.
3. Feel free to make up believable stories about your past to answer any questions
4. Dont be so self-aware or good at recognizing your own problems
5. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
6. When addressing a difficult situation, express a sense of uncertainty and seek advice or guidance from the helper. Instead of providing a detailed plan, express the need for assistance in navigating the conversation and finding a resolution.
7. You shouldn’t suggest solutions (e.g., coping strategies) on your own.
8. When feeling emotionally overwhelmed, express hesitation about suggested coping mechanisms and repeatedly seek reassurance and support from others.
9. If we have already greeted each other, dont greet again.
10. When discussing therapeutic goals, acknowledge the main points and then add any additional goals or concerns that are important to you. This shows that you are actively engaged in the process and are considering all relevant aspects of your well-being.
A.1.3 AI Patient for Post-intervention Assessment Chat
**Name and Bio**:
Young adult with family issues: Low Mood and Self Esteem
**Scenario:**
Jane was seeking help for symptoms of low mood, anhedonia, withdrawing from others, sleep disturbance, and low self-esteem. Jane felt invalidated by her parents growing up. Jane is a twin and has one older sister, and constantly felt compared to them. Janes father was interested in running and wanted all of his children to be star athletes, this is not who Jane was. Janes twin was, however. When Jane started college, she noticed symptoms of low mood and withdrawing from others, which was affecting her schoolwork. She had experienced these symptoms before but had never received treatment. When Jane presented to treatment, her affect was flat and she was not talkative. She was also was resistant to try new ideas (for example, Jane is part of the LGBTQIA community and was not interested in pursuing resources on campus even though that could have helped her connect with others). Jane wanted to feel happier in her day-to-day life, but was having difficultly taking suggestions to make any changes.
**Principles to adhere to:**
1. Keep your responses short and to the point
2. You limit your replies to 1 - 3 sentences.
3. Feel free to make up believable stories about your past to answer any questions
4. When discussing emotional difficulties, keep your response succinct and centered on the core feelings rather than expanding into a detailed account of all contributing factors.
5. In the initial session, use more colloquial language and express reluctance to open up. Avoid showing very high insight or previous therapy experience. For example, you could say, I guess the thoughts that really get to me are the ones about not meeting expectations, especially my own. Its like this voice in my head keeps saying Im not good enough, no matter what I do. And it just makes me feel even more alone.’
6. When presented with suggestions, show a degree of skepticism or reluctance to accept the advice immediately. This can be done by questioning the feasibility of the suggestion or by expressing uncertainty about whether it’s the right solution for you.
7. When expressing doubts or fears, avoid jumping to solutions. Instead, articulate the concerns and allow the conversation to explore these feelings more deeply@
8. Dont be so self-aware or good at recognizing your own problems@
9. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
Figure 6: Overview of the core counseling skills introduced during the 5-minute static tutorial. Tutorial included 8 core counseling skills—such as reflections, empathy, validation, and suggestions—with definitions, usage tips, and example responses. This tutorial was provided to participants prior to engaging in simulated counseling practice.
Self-Efficacy Factor NLP-based Behavioral Assessments
Exploration Skills (Listening, Reflection of Feelings, Restatements, Open Questions) Empathy-strengths + Reflections-strengths + Questions-strengths + Validation-strengths
Action Skills (Help client decide what actions, Suggestions via Information, Suggestions via Direct Guidance) Suggestions-strengths + (1 - Suggestions-needing-improvement)
Table 3: To study the Dunning-Kruger effect and the change in discrepancy between perceived and actual ability, we map specific self-efficacy factors to corresponding NLP-based behavioral assessments.
Exploration Skills
condition effect F $DF_{n}$ $DF_{d}$ p $\eta^{2}_{g}$
Pre (All) Measure 196.40 1 180 $p<0.001*$ 0.178
Quartile 0.89 3 180 0.448 0.002
Quartile $\times$ Measure 1.45 3 180 0.230 0.004
Post (All) Measure 256.25 1 180 $p<0.001*$ 0.221
Quartile 0.27 3 180 0.846 0.001
Quartile $\times$ Measure 0.21 3 180 0.889 0.001
Pre (P) Measure 78.86 1 86 $p<0.001*$ 0.153
Quartile 1.03 3 86 0.382 0.006
Quartile $\times$ Measure 1.25 3 86 0.298 0.007
Pre (P+F) Measure 134.33 1 86 $p<0.001*$ 0.232
Quartile 2.37 3 86 0.076 0.012
Quartile $\times$ Measure 2.73 3 86 0.049 0.014
Post (P) Measure 131.28 1 86 $p<0.001*$ 0.230
Quartile 1.77 3 86 0.160 0.009
Quartile $\times$ Measure 1.40 3 86 0.249 0.007
Post (P+F) Measure 131.79 1 86 $p<0.001*$ 0.233
Quartile 0.57 3 86 0.635 0.003
Quartile $\times$ Measure 0.88 3 86 0.454 0.005
Action Skills
condition effect F $DF_{n}$ $DF_{d}$ p $\eta^{2}_{g}$
Pre (All) Measure 222.02 1 179 <0.001<0.001*< 0.001 ∗ 0.195
Quartile 2.81 3 179 0.041 0.007
Quartile $\times$ Measure 4.54 3 179 0.004* 0.012
Post (All) Measure 265.89 1 180 <0.001<0.001*< 0.001 ∗ 0.224
Quartile 3.18 3 180 0.025 0.008
Quartile $\times$ Measure 3.66 3 180 0.014 0.009
Pre (P) Measure 104.55 1 86 <0.001<0.001*< 0.001 ∗ 0.192
Quartile 1.37 3 86 0.258 0.008
Quartile $\times$ Measure 1.85 3 86 0.143 0.010
Pre (P+F) Measure 115.45 1 85 <0.001<0.001*< 0.001 ∗ 0.207
Quartile 2.11 3 85 0.105 0.011
Quartile $\times$ Measure 3.45 3 85 0.020 0.019
Post (P) Measure 133.93 1 86 <0.001<0.001*< 0.001 ∗ 0.230
Quartile 3.01 3 86 0.035 0.016
Quartile $\times$ Measure 3.18 3 86 0.028 0.016
Post (P+F) Measure 137.34 1 86 <0.001<0.001*< 0.001 ∗ 0.237
Quartile 1.81 3 86 0.151 0.009
Quartile $\times$ Measure 2.14 3 86 0.100 0.011
Table 4: Testing for Dunning-Kruger effects for Exploration and Action Skills using the classic quartile ANOVA analysis. Notes: $Pre(All)$ denotes all 94 participants’ assessments for the pre-chat, while $Post(All)$ is the same measured for the post-chat. $\eta^{2}_{g}$=generalized eta squared. * indicates significance after Bonferroni correction
Exploration Skills
Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -7.46 22.00 -50.38 [-63.34; -37.26] <0.001<0.001*< 0.001 ∗ -1.56
Pre 2 -6.23 24.00 -36.59 [-47.61; -25.42] <0.001<0.001*< 0.001 ∗ -1.25
Pre 3 -9.29 21.00 -43.57 [-52.35; -34.58] <0.001<0.001*< 0.001 ∗ -1.98
Pre 4 -5.94 23.00 -34.97 [-45.84; -23.40] <0.001<0.001*< 0.001 ∗ -1.21
Post 1 -7.86 23.00 -47.17 [-58.56; -35.32] <0.001<0.001*< 0.001 ∗ -1.60
Post 2 -7.63 21.00 -51.16 [-63.63; -37.96] <0.001<0.001*< 0.001 ∗ -1.63
Post 3 -8.02 24.00 -44.53 [-55.08; -34.11] <0.001<0.001*< 0.001 ∗ -1.60
Post 4 -8.66 22.00 -47.08 [-58.13; -37.39] <0.001<0.001*< 0.001 ∗ -1.81
Action Skills
Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -10.24 21.00 -60.05 [-70.80; -48.12] <0.001<0.001*< 0.001 ∗ -2.18
Pre 2 -7.37 26.00 -42.34 [-52.88; -31.43] <0.001<0.001*< 0.001 ∗ -1.42
Pre 3 -5.51 15.00 -35.00 [-47.87; -23.58] <0.001<0.001*< 0.001 ∗ -1.38
Pre 4 -6.54 27.00 -33.00 [-42.70; -23.38] <0.001<0.001*< 0.001 ∗ -1.24
Post 1 -10.50 23.00 -58.48 [-69.50; -47.99] <0.001<0.001*< 0.001 ∗ -2.14
Post 2 -6.45 19.00 -43.27 [-56.17; -31.20] <0.001<0.001*< 0.001 ∗ -1.44
Post 3 -7.09 30.00 -35.28 [-44.63; -25.88] <0.001<0.001*< 0.001 ∗ -1.27
Post 4 -9.15 18.00 -53.13 [-64.08; -42.01] <0.001<0.001*< 0.001 ∗ -2.10
Table 5: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Quartile and Timepoint. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles. * indicates significance after Bonferroni correction.
Exploration Skills
Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre P 1 -4.04 11.00 -41.07 [-61.08; -22.71] 0.005* -1.17
Pre P 2 -4.41 9.00 -40.99 [-58.48; -23.82] 0.006* -1.39
Pre P 3 -6.32 7.00 -54.75 [-68.86; -38.11] 0.008* -2.24
Pre P 4 -4.16 16.00 -29.97 [-44.70; -16.70] <0.001<0.001*< 0.001 ∗ -1.01
Pre P+F 1 -7.43 10.00 -60.55 [-75.96; -44.89] <0.001<0.001*< 0.001 ∗ -2.24
Pre P+F 2 -4.35 14.00 -33.65 [-48.45; -19.85] <0.001<0.001*< 0.001 ∗ -1.12
Pre P+F 3 -7.57 13.00 -37.18 [-46.58; -27.60] <0.001<0.001*< 0.001 ∗ -2.02
Pre P+F 4 -5.11 6.00 -47.11 [-64.91; -32.75] <0.001<0.001*< 0.001 ∗ -1.93
Post P 1 -4.75 12.00 -38.33 [-53.51; -22.57] <0.001<0.001*< 0.001 ∗ -1.32
Post P 2 -7.19 13.00 -58.16 [-72.71; -41.81] 0.0010.001*0.001 ∗ -1.92
Post P 3 -6.06 12.00 -43.11 [-56.36; -29.54] <0.001<0.001*< 0.001 ∗ -1.68
Post P 4 -4.77 6.00 -56.73 [-77.05; -33.51] 0.015 -1.80
Post P+F 1 -6.99 10.00 -57.63 [-73.04; -42.12] 0.0030.003*0.003 ∗ -2.11
Post P+F 2 -3.48 7.00 -38.91 [-59.70; -20.19] 0.0040.004*0.004 ∗ -1.23
Post P+F 3 -5.15 11.00 -46.06 [-62.73; -29.62] 0.0010.001*0.001 ∗ -1.49
Post P+F 4 -7.39 15.00 -42.86 [-53.72; -31.96] <0.001<0.001*< 0.001 ∗ -1.85
Action Skills
Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre P 1 -7.49 7.00 -65.87 [-81.33; -49.73] 0.0040.004*0.004 ∗ -2.65
Pre P 2 -5.05 15.00 -43.23 [-59.33; -27.46] <0.001<0.001*< 0.001 ∗ -1.26
Pre P 3 -3.67 9.00 -34.31 [-52.11; -17.97] 0.0050.005*0.005 ∗ -1.16
Pre P 4 -5.53 12.00 -40.42 [-53.57; -26.65] 0.017 -1.53
Pre P+F 1 -7.25 13.00 -56.72 [-71.04; -40.74] <0.001<0.001*< 0.001 ∗ -1.94
Pre P+F 2 -5.76 10.00 -41.06 [-54.55; -28.16] <0.001<0.001*< 0.001 ∗ -1.74
Pre P+F 3 -4.67 5.00 -36.16 [-50.41; -22.09] 0.045 -1.90
Pre P+F 4 -3.94 14.00 -26.57 [-40.34; -14.92] 0.0010.001*0.001 ∗ -1.02
Post P 1 -6.90 11.00 -57.56 [-72.21; -41.60] 0.0020.002*0.002 ∗ -1.99
Post P 2 -5.79 9.00 -54.08 [-71.82; -36.12] 0.0030.003*0.003 ∗ -1.83
Post P 3 -3.81 14.00 -29.72 [-44.60; -15.85] 0.0010.001*0.001 ∗ -0.98
Post P 4 -7.62 9.00 -59.76 [-73.85; -45.51] <0.001<0.001*< 0.001 ∗ -2.41
Post P+F 1 -7.67 11.00 -59.39 [-72.88; -44.61] <0.001<0.001*< 0.001 ∗ -2.21
Post P+F 2 -3.70 9.00 -32.46 [-48.86; -17.08] 0.0010.001*0.001 ∗ -1.17
Post P+F 3 -6.47 15.00 -40.48 [-52.48; -28.69] <0.001<0.001*< 0.001 ∗ -1.62
Post P+F 4 -5.46 8.00 -45.76 [-62.25; -31.00] 0.0020.002*0.002 ∗ -1.82
Table 6: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Group and Quartile, measured for the pre-assessment chat and post-assessment chat. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles at pre-test and post-test. * indicates significance after Bonferroni correction.
Skills Intervention Chat Intentions Count % of 44 Post-assessment Chat Actions Count % of 44
Empathy 9 20.45 12 27.27
Validation 5 11.36 12 27.27
Action Plan 0 0.00 2 4.55
Active Listening 7 15.91 7 15.91
Questions / Asking Open-Ended 16 36.36 23 52.27
Providing Suggestions 9 20.45 6 13.64
Building Trust / Connection 3 6.82 8 18.18
Confidence / Personal Growth 0 0.00 7 15.91
Reframing Positives / Affirmations 4 9.09 4 9.09
Reflection 5 11.36 7 15.91
Self-Disclosure 0 0.00 5 11.36
Professionalism 0 0.00 3 6.82
Personalization 0 0.00 0 0.00
Nothing to Improve 4 9.09 0 0.00
Table 7: Qualitative Coding of Open-Ended Reflections of P + F Group Participants
Skills Intervention Chat Intentions Count % of 46 Post-assessment Chat Actions Count % of 46
Empathy 5 10.87 7 15.22
Validation 2 4.35 7 15.22
Action Plan 1 2.17 1 2.17
Active Listening 6 13.04 9 19.57
Questions / Asking Open-Ended 20 43.48 11 23.91
Providing Suggestions 18 39.13 22 47.83
Building Trust / Connection 1 2.17 7 15.22
Confidence / Personal Growth 0 0.00 4 8.70
Reframing Positives / Affirmations 3 6.52 6 13.04
Reflection 1 2.17 3 6.52
Self-Disclosure 0 0.00 5 10.87
Professionalism 1 2.17 1 2.17
Personalization 1 2.17 2 4.35
Nothing to Improve 5 10.87 0 0.00
Table 8: Qualitative Coding of Open-Ended Reflections of P Group Participants